OpenStudy (anonymous):

URGENT!! HELP ME PLEASE! given dy/dx = a(e^-x)+2 and that when x=0, dy/dx =5 and y=1, then when x=2, what is y equals to?

zepdrix (zepdrix):

Do you know how to anti-differentiate? Or are expected to solve this using a different method?

OpenStudy (anonymous):

you have to use anti diff. i know how to anti differentiate but i cant get the answer :(

zepdrix (zepdrix):

Let's start by solving for the missing A value c: Plugging in the given values will give us,\[\large \frac{dy}{dx}=ae^{-x}+2 \qquad \rightarrow \qquad 5=ae^0+2 \qquad \rightarrow \qquad 3=a\]Were you able to get that part? c:

OpenStudy (anonymous):

yup!!

zepdrix (zepdrix):

\[\large \frac{dy}{dx}=3e^{-x}+2\]We'll rewrite the dx on the other side,\[\large dy=3e^{-x}+2 \; dx\] Integrating gives us,\[\large \int\limits dy=\int\limits 3e^{-x}+2 \; dx \qquad \rightarrow \qquad y=-3e^{-x}+2x+C\]

OpenStudy (anonymous):

there i found that i cant get C :(

zepdrix (zepdrix):

Any confusion? If you're just being introduced to "anti-differentiation", then the integral sign might be confusing. I forgot about that maybe :D

zepdrix (zepdrix):

Ok let's seeeeee

zepdrix (zepdrix):

So they gave us an initial condition to use. When x=0, y=1. Plugging those in gives us,\[\large 1=-3e^0+2\cdot 0 + C\] Hmm I'm not sure why they gave us an initial condition containing information for the derivative and anti-derivative. That's kinda strange... I mean I guess it was so we could find A, then C later on. It's just weird :3 Hopefully I didn't make a silly error in there anywhere.

OpenStudy (anonymous):

it is 8 :(

OpenStudy (anonymous):

err..

zepdrix (zepdrix):

Oh that 2 was being multiplied by 0, woops my bad. So C=4 I think? :o

zepdrix (zepdrix):

We'll plug our new-found C value into our equation,\[\large y=-3e^{-x}+2x+4\]

zepdrix (zepdrix):

They give us another x value, and ask us to find the corresponding y value. So when x=2,\[\large y=-3e^{-2}+2\cdot 2+4\]